DECELERATION, TRAPPING AND ACCUMULATION OF NH MOLECULES

\underline{STEVEN HOEKSTRA}, JOOP J. GILIJAMSE, SEBASTIAAN Y.T. VAN DE MEERAKKER and GERARD MEIJER, Fritz-Haber-Institut der Max-Planck-Gesellschaft, Berlin, Germany.

We report on the Stark-deceleration and electrostatic trapping of metastable NH molecules. Furthermore the progress towards higher densities of cold neutral molecules by accumulation of multiple Stark-decelerated packets of NH molecules in a magnetic trap will be presented. NH molecules in the long-lived metastable a^{1}Δ(v = 0, J = 2) state are ideally suited for Stark deceleration experiments because of their relatively large Stark shift and low mass. The metastable molecules (τ > 2.7s) are produced in a supersonic expansion with a velocity of ∼ 450 m/s, and are decelerated to a standstill by a 108-stage decelerator. Subsequently the metastable molecules are trapped electrostatically, with a temperature of about 50 − 100 mK, a density of ∼ 10^{6} cm^{−3} and a 1/e trapping lifetime of 1.4 s. Following the deceleration and trapping, the metastable NH molecules are detected by the excitation of a spin-forbidden transition, resulting in spontaneous decay to the electronic ground state (X^{3}Σ^{−}). The electronic ground state has a negligible Stark shift, but can be trapped magnetically. The first experiments on the accumulation of ground state NH molecules in a magnetic trap will be presented.

[1] S. Hoekstra et al., Electrostatic trapping of metastable NH molecules, Phys. Rev. A. 76 063408 (2007)